40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2143     Entry time: Mon Oct 26 17:45:34 2009
Author: Jenne 
Type: Update 
Category: Adaptive Filtering 
Subject: New changes to the OAF fe code 

[Alex, Jenne, Sanjit]

Alex came to the 40m today, and did several awesome things in OAF-land.

We discovered that there is, in fact, an ADC board connected to the ASS machine.  The tricky bit is that it only has a ribbon cable connector, so before we can use this ADC, we need to figure out how to make a breakout board/cable/something to connect the seismometer/accelerometer/microphone BNCs to this little board.  This is the same little board that connects the timing slave to the ASS machine.  For good or for ill, the timing slave is connected to this board via clip-doodles.  Potentially we can connect an ADC tester board to this board, and go from seismometer BNCs to clipdoodles to the tester board, but I'm not in love with the idea of utilizing clipdoodles as a semi-permanent solution until the upgrade.  I emailed Ben to see if he has a better idea, or (better yet) some spare hardware now that's the same as we'll use after the upgrade.  If we can use this ADC, it may solve our timing problem which is caused by the 110B ADC used by the PEM computer. Alex showed Sanjit and I how to connect the ASS's ADC card to the simulink diagram, when we're ready for that.

We also poked around in the code, and it seems that we can now save and restore OAF coefficients at will.  I added buttons to the OAF (ASS) screen, and Alex made it so the OAF coefficients are saved in RFM shared memory whenever you click the "save coeffs" button, and are restored when you click the "restore coeffs" button.  The buttons are the same as the 'Reset' button which has been there for a long time, so they seem to maybe have a similar problem in that you have to hold the button for a while in order for the code to realize that the button has been depressed.  We couldn't fix this easily, because it looks like our SimuLink cds stuff is a little out of date.  Some day (before/when Joe and Peter make new screens for the new 40m), we need to update these things.  Alex was concerned that it might take a while to do this, if the update broke some of the blocks that we're currently using.  Also, Sanjit and I now need to check that the coefficient-saving is going as planned.  When I have DTT open, and the OAF running, I see a certain shape to the signal which is sent to MC1 to correct for the seismic motion.  This shape includes at least several peaks at resonant frequencies that exist in our stacks/suspensions.  I can then save the coefficients, reset the active filter, and then restore the coefficients.  When I do this while watching DTT, it seems as though the general shape of the filter is restored, but none of the detailed features are.  The reason for this is still under investigation. 

The code-modifications involved a few iterations of 'remaking the ass'.

ELOG V3.1.3-